OpenStudy (anonymous):

@katalina13 come in here and i will help you with it all

OpenStudy (me110000):

ok here

OpenStudy (anonymous):

ya it is for a another person but if you need help with anything i am willing to help

OpenStudy (anonymous):

Mrs. miller sells a house 179,000. If she earnsa commition of 6 % how much does she earn. Write a paportion and show your work

OpenStudy (me110000):

ok Srry My name is katalina

OpenStudy (anonymous):

its a pretty name

OpenStudy (anonymous):

you are ok

OpenStudy (anonymous):

katalina13 what is your real name

OpenStudy (anonymous):

katelynn

OpenStudy (anonymous):

that is what my best friends baby's name is

OpenStudy (me110000):

ok @katalina13 Check this link out: http://openstudy.com/study#/updates/50ad5e95e4b09749ccabbbe3

OpenStudy (anonymous):

6% = 6/100 ...can you write what 6% of 179000

OpenStudy (anonymous):

part/whole = percent/100 x/179000 = 6/100 x/179000 = 0.06 x = 0.06*179000 x = 10740 So you are 100% correct

OpenStudy (anonymous):

i got the answer

OpenStudy (anonymous):

thanks